OpenStudy (anonymous):

What is the radius of a circle with the equation x^2+y^2+2x+4y-9=0?

OpenStudy (anonymous):

You have to complete the square for both x^2 + 2x and y^2 +4y

OpenStudy (anonymous):

Okay, I just wasn't too sure about this one, but that's what I thought. Thank you.

OpenStudy (anonymous):

Wait, do I just plug those values in after that?

OpenStudy (anonymous):

x^2 + 2x = (x+1)^2 - 1 y^2 +4y = (y+2)^2 - 4 (x+1)^2 - 1 + (y+2)^2 - 4 - 9 = 0 Move the constants to the right and take the square root